OpenStudy (anonymous):

. In the following equation, identify the x-intercepts in the graph y = 3x2 + 19x – 40

OpenStudy (***[isuru]***):

x intercepts mean the value of x when y = o

OpenStudy (***[isuru]***):

take y = 3x2 + 19x – 40 and when y = 0 0 = 3x2 + 19x – 40 now solve this for x...that the answer.. may i help to continue or can u continue this from here ?

OpenStudy (anonymous):

can you keep going?

OpenStudy (***[isuru]***):

u can use the formula to find the x value .. when \[ax^{2} + bx + c = 0\] \[x = \frac{ -b \pm \sqrt{b^{2} - 4ac} }{ 2a }\]

OpenStudy (***[isuru]***):

in the given expression 3x^2 + 19x -40 = 0 a = 3 b = 19 c = (-40) now plug those values to the formula

OpenStudy (***[isuru]***):

\[x = \frac{ -19 \pm \sqrt{19^{2} - 4[3\times(-40) }] }{ 6 }\]

OpenStudy (***[isuru]***):

\[x = \frac{ -19 \pm \sqrt{361 + 480} }{ 6}\]